# Optomechanical Systems with Linear and Quadratic Position Couplings: Dynamics and Optimal Estimation

### v1

@@ Abstract
We study the dynamics of an optomechanical system consisting of a single-mode optical field coupled to a mechanical oscillator, where the nonlinear interaction includes both linear and quadratic terms in the oscillator’s position. We present a full analytical solution to this quantum mechanical Hamiltonian problem by employing the formalism of two-phonon coherent states. Quantum estimation theory is applied to the resulting state of the optical field, with a focus on evaluating the classical and quantum Fisher information for estimating the strength of the quadratic coupling. Our estimation scheme considers both standard and balanced homodyne photodetection, assuming an initial optical state prepared as a superposition of vacuum and single-photon states. We show that balanced homodyne detection can saturate the quantum Fisher information, thus reaching the ultimate precision bound for estimating the quadratic coupling. Additionally, we investigate the effect of thermal noise on the quantum Fisher information in a realistic experimental context.
@@ I Introduction
Optomechanical systems [1] are at the center of a technological revolution [2] that is based on the interaction between light and matter to study fundamental physics [3] and produce innovative devices, spanning from signal processors [4] to transductors [5] and quantum amplifiers [6, 7]. Achieving these goals requires a thorough understanding of the optomechanical system, which can be attained through precise measurements of its parameters. Numerous studies [8, 9, 10, 11, 12, 13] have focused on understanding how to estimate the parameters of an optical cavity coupled to a mechanical oscillator via radiation pressure.
@@
In constructing effective estimation frameworks, the first essential step is to define a statistical model for the data [14, 15]. In quantum mechanics, this requirement translates to obtaining an exact solution for the time evolution of the quantum state. Formally, this yields a family of quantum states parameterized by the unknown variables to be estimated. However, due to the complexity introduced by the high-dimensional parameter spaces commonly encountered in optomechanical systems, it is often necessary to introduce simplifications to the physical model. These approximations facilitate the formulation of an minimal model that remains tractable while still capturing the essential physics. A detailed derivation of the optomechanical Hamiltonian is presented in the foundational works [16, 17]. Two principal approximations are typically employed to simplify this Hamiltonian. First, if the mechanical oscillator evolves adiabatically slowly compared to the frequency separation between optical modes, intermodal photon scattering can be neglected. Under this condition, the system can be effectively described using a single optical cavity mode. Second, in the linear approximation, the radiation–mechanical interaction is simplified by expanding the coupling Hamiltonian to first order in the mechanical position operator. The first objective of this paper is to construct a statistical model by deriving the exact time-evolving quantum state governed by the optomechanical Hamiltonian, under the adiabatic approximation while extending the standard linear interaction to include second-order (quadratic) terms in the mechanical displacement.
@@
The Hamiltonian under investigation has been previously discussed in comprehensive reviews [1] and has found applications in various contexts, including quantum nondemolition measurements of the phonon number of the mechanical mode [18], as well as in optomechanically induced parametric oscillations [19]. Mathematical treatments often focus on the corresponding Heisenberg equations of motion, their expectation values, and employ approximations to derive a finite set of coupled differential equations for the system dynamics [20]. Our focus is on obtaining an exact solution for the time-evolving quantum state, which, to our knowledge, has not been previously derived. We employ the formalism of two-photon coherent states [21], adapted here to describe the mechanical oscillator, allowing us to call them as two-phonon coherent states. By working in the Fock basis of the single-mode optical field, the Hamiltonian becomes block diagonal, with each block corresponding to a fixed photon number. The dynamics within each block are then governed by the evolution of the photon number-related two-phonon coherent states. The result describes the joint quantum state of the optical field and the mechanical oscillator. Since, in practice, measurements are typically performed on the optical field, we trace out the mechanical degrees of freedom. We therefore focus on the resulting reduced state of the optical field. This reduced optical state serves as the basis for the quantum estimation procedure.
@@
Here, we adopt a frequentist approach and focus on the lower bounds of the variance for any unbiased estimator. Even without explicitly constructing the estimators, these bounds serve as benchmark values for assessing the performance of those used in experiments. In this context, we analyze the quantum Fisher information (QFI), which represents the lower bound in the Cramér Rao inequality [22, 23], for the optical field state as a function of the unknown strength of the optomechanical coupling, which depends quadratically on the mechanical position operator. This is compared with the classical Fisher information (CFI), obtained for standard and balanced-homodyne photodetection measurement schemes. The QFI serves as a measure of the sensitivity of the optical state to changes in the quadratic optomechanical coupling constant; in other words, it quantifies how much information the quantum state contains about this parameter. On the other hand, the CFI assesses the potential to evaluate this sensitivity using a set of classical measurements. For simplicity, we assume that the mechanical oscillator is initially in a thermal state, while the optical field is prepared in a superposition of the vacuum and single-photon Fock states. It is important to note that the cavity is assumed to be lossless, and its quantum state can be accessed without inducing any disturbance.
@@ II Model
where $\hat{x}$ is the position operator of the mechanical system, $\hat{p}$ is its momentum operator, $m$ is its mass, and $\Omega$ is the mechanical frequency of motion. The operators $\hat{a},\hat{a}^{\dagger}$ are the photonic ladder operators with the commutation relation $[\hat{a},\hat{a}^{\dagger}]=1$, and $\omega(x)$ is the resonant optical frequency of the cavity, which depends on the position $x$ of the mechanical oscillator. If we expand the optical frequency $\omega(x)$ around the equilibrium value $x_{0}$, we obtain
@@
where $\omega^{(n)}(x_{0})$ denotes the $n$th derivative of $\omega(x)$ evaluated at the point $x_{0}$. Here, $g_{1}$ is the linear, and $g_{2}$ is the quadratic optomechanical coupling constant. The parameter $\omega_{c}$ is identified as the frequency of the single mode of the radiation field. By considering the position as an operator, we can write the optomechanical Hamiltonian as
@@
It is worth mentioning that the single-mode approach requires the motion of the mechanical oscillator to be adiabatically slow, so the model can ignore the scattering of photons from the mode to other cavity modes [16].
@@
where $\hat{I}$ is the identity operator on the Hilbert space of the mechanical oscillator. By setting $\Omega_{n}=\sqrt{\Omega^{2}+\frac{2\hbar g_{2}n}{m}}$, we can define the ladder mechanical operators $\hat{b}_{n},\hat{b}^{\dagger}_{n}$ as
@@
where $\omega_{n}=n\omega_{c}+\Omega_{n}/2$ and $g_{1,n}=ng_{1}\sqrt{\frac{\hbar}{2m\Omega_{n}}}$. By applying the Baker-Campbell-Hausdorff formula, we get [24]
@@
Given an initial state $\rho(0)$, the time evolution of the system is given by the von Neumann equation
$\hat{\rho}(t)=e^{-i\hat{H}^{\text{quad}}t/\hbar}\hat{\rho}(0)e^{i\hat{H}^{\text{quad}}t/\hbar}.$ (13)
@@
$\hat{\rho}(0)=\hat{\sigma}(0)\otimes\mathfrak{\hat{m}}(0),$ (14)
where $\hat{\sigma}(0)$ and $\mathfrak{m}$ are the optical and the mechanical density operator, respectively.
@@
where $d^{2}\alpha=d\mathrm{Re}\{\alpha\}\,d\mathrm{Im}\{\alpha\}$, $\lvert{\alpha}\rangle_{0}$ is a coherent state, and $P(\alpha,\alpha^{*})$ is a Glauber-Sudarshan quasidistribution [25, 26]. Due to the fact that $\hat{H}^{\text{quad}}$ is block diagonal in the photonic Fock basis, we choose
$\hat{\sigma}(0)=\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert,$ (16)
@@
$\mathrm{Tr}\{\hat{\sigma}(0)\}=1\quad\text{and}\quad\hat{\sigma}(0)\geq 0,$ (17)
i.e., $\hat{\sigma}(0)$ is a positive semidefinite operator with trace one.
@@
$\hat{\rho}(t)$ $=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\hat{H}_{n}t/\hbar}\hat{\mathfrak{m}}(0)e^{i\hat{H}_{m}t/\hbar}$ (22)
@@
$\hat{\rho}(t)$ $=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\left[\Phi_{n}(t)-\Phi_{m}(t)\right]}$
@@
The remainder of the calculation will focus solely on the optical density operator, $\hat{\sigma}(t)$, at time $t$. This approach is justified by the fact that, in most practical scenarios, the optical field is the only component of the system that is experimentally accessible. The operator $\hat{\sigma}(t)$ is obtained by performing a partial trace over the mechanical degrees of freedom, and it can be expressed as
$\hat{\sigma}(t)=\mathrm{Tr}_{\text{mech}}\{\hat{\rho}(t)\}=\sum^{\infty}_{n,m=0}a_{n,m}(t)|n\rangle_{c}\langle m|,$ (28)
@@
The complete expressions for $A$, $b$ and $c$ are provided in Appendix B, all derived from the overlaps of two-phonon coherent states [29], as explicitly given in Eq. (LABEL:eq:A1). The coefficients $a_{n,m}(t)$ of the optical density operator provide a complete description of the interaction between the single mode of the radiation field and the single vibrational mode of the mechanical oscillator, under the assumption that all losses and sources of decoherence are neglected. We now use these results to estimate the strength of the quadratic optomechanical coupling.
@@ III Quantum and Classical Fisher information
$\text{Var}_{g}\geq\frac{1}{N_{\text{meas}}\text{CFI}}\geq\frac{1}{N_{\text{meas}}\text{QFI}}.$ (35)
@@
In quantum mechanics, the measurement procedure is described by a Positive Operator-Valued Measure (POVM), a collection of operators $\{\Pi_{i}\}_{i}$, such that any $\Pi_{i}\geq 0$ and $\sum_{i}\Pi_{i}=\mathds{1}$. The CFI of the state $\rho$ and the measurement $\Pi$ can be calculated as
$\text{CFI}=\sum_{i}P_{i}[g]\bigg{(}\frac{\partial\ln P_{i}[g]}{\partial g}\bigg{)}^{2},$ (36)
where $P_{i}[g]=\text{Tr}[\rho_{g}\Pi_{i}]$ is the probability of getting the outcome $i$ when the true value of the parameter is $g$.
@@
$\text{QFI}=\text{Tr}[\rho_{g}L_{g}^{2}]$ (37)
which is written in terms of the symmetric logarithmic derivative (SLD) $L_{g}$ [32], which satisfies the relation
$\frac{\partial\rho_{g}}{\partial g}=\frac{1}{2}(L_{g}\rho+\rho L_{g}).$ (38)
@@
[…]
We mention here that when the density operator has the peculiar property that $\sigma^{2}=\sigma-h\mathbf{I}$, with $h\in\mathbb{R}$, which always happens when $\sigma$ describes a two-level system, the SLD assumes the simplified form [33]
$L_{g}=2\partial_{\theta}\rho-\frac{1}{2}\partial_{\theta}\mathcal{P}\rho^{-1},$ (40)
where $\mathcal{P}=\text{Tr}[\rho^{2}]$ is the purity of the system and is related to $h$ by
$\mathcal{P}=2h+1,$
which can be used to ease the computation of the QFI. In the following we are going to apply Eq. (40) to calculate the QFI for the case when the cavity is populated by a superposition of zero-photon and one-photon states.
In the case of a $2\times 2$ system we can define an analogue of the photodetection and the balanced homodyne detection (BHD) measurements in terms of their projective value measures (PVM). For the photodetection, the corresponding PVM is $\{\Pi_{\text{ph}}\}$, with
$\Pi_{\text{ph}}(0)=|0\rangle\langle 0|$ (41)
$\Pi_{\text{ph}}(1)=|1\rangle\langle 1|.$
The BHD [34] enables the measurement of the observable
$\hat{X}_{\phi}=\frac{\hat{a}e^{i\phi}+\hat{a}^{\dagger}e^{-i\phi}}{2}.$ (42)
which, we rewrite to account for the two-level system scenario that we are considering by substituting the operators $\hat{a},\hat{a}^{\dagger}$ with the $2\times 2$ counterparts $\sigma^{-},\sigma^{+}$ respectively. The corresponding PVM $\{\Pi_{X_{\phi}}\}$ is given by the projectors on the two eigenstates $|e_{0,1}\rangle$ of $X_{\phi}$,
$\Pi_{X_{\phi}}(k)=|e_{k}\rangle\langle e_{k}|,\quad k=1,2.$ (43)

### v2

@@ Abstract
We study the dynamics of an optomechanical system consisting of a single-mode optical field coupled to a mechanical oscillator, where the nonlinear interaction includes both linear and quadratic terms in the oscillator’s position. We present an analytical solution to this quantum mechanical Hamiltonian problem by employing the formalism of two-phonon coherent states. Quantum estimation theory is applied to the resulting state of the optical field, with a focus on evaluating the quantum Fisher information with respect to the strength of the quadratic coupling. Our estimation scheme employs balanced homodyne photodetection and demonstrates that the corresponding classical Fisher information can reach the quantum Fisher information limit, with the phase of the local coherent oscillator playing a crucial role.
@@ I Introduction
Optomechanical systems [2] are at the center of a technological revolution [3] that is based on the interaction between light and matter to study fundamental physics [26] and produce innovative devices, spanning from signal processors [1] to transductors [8] and quantum amplifiers [23, 32]. Achieving these goals requires a thorough understanding of the optomechanical system, which can be attained through precise measurements of its parameters. Numerous studies [5, 30, 34, 6, 28, 31, 25, 24] have focused on understanding how to estimate the parameters of an optical cavity coupled to a mechanical oscillator via radiation pressure.
@@
In constructing effective estimation frameworks, the first essential step is to define a statistical model for the data [40, 15]. In quantum mechanics, this requirement translates to obtaining an exact solution for the time evolution of the quantum state. Formally, this yields a family of quantum states parameterized by the unknown variables to be estimated. However, due to the complexity introduced by the high-dimensional parameter spaces commonly encountered in optomechanical systems, it is often necessary to introduce simplifications to the physical model. These approximations facilitate the formulation of a minimal model that remains tractable while still capturing the essential physics. A detailed derivation of the optomechanical Hamiltonian is presented in the foundational works [17, 18]. Two principal approximations are typically employed to simplify this Hamiltonian. First, if the mechanical oscillator evolves adiabatically slowly compared to the frequency separation between optical modes, intermodal photon scattering can be neglected. Under this condition, the system can be effectively described using a single optical field mode. Second, in the linear approximation, the radiation–mechanical interaction is simplified by expanding the coupling Hamiltonian to first order in the mechanical position operator. The first objective of this paper is to construct a statistical model by deriving the exact time-evolving quantum state governed by the optomechanical Hamiltonian, under the adiabatic approximation while extending the standard linear interaction to include second-order (quadratic) terms in the mechanical displacement.
@@
The Hamiltonian under investigation has been previously discussed in comprehensive reviews [2]. It has found applications in various contexts, including quantum nondemolition measurements of the phonon number of the mechanical mode [9], as well as in optomechanically induced parametric oscillations [27]. Mathematical analyses of such systems often concentrate on the corresponding Heisenberg equations of motion and their expectation values. In many cases, these approaches rely on suitable approximations to obtain a finite set of coupled differential equations describing the system dynamics [16]. An alternative strategy is to confine the dynamics to Gaussian states, which enables the use of general closed form expressions in quantum parameter estimation studies [29]. The aforementioned approaches can even be extended to include dissipative processes within the Hamiltonian framework. In contrast, the present work is concerned with obtaining an exact analytical solution for a general time-dependent quantum state, without imposing Gaussian or perturbative approximations. To the best of our knowledge, such an exact formulation has not been reported previously. We employ the formalism of two-photon coherent states [42], adapted here to describe the mechanical oscillator. The obtained solutions describe the joint quantum state of the optical field and the mechanical oscillator. Since, in practice, measurements are typically performed on the optical field, we trace out the mechanical degrees of freedom. We provide a comprehensive analysis on the resulting state and its dynamics.
@@
In the quantum parameter estimation analysis, we employ a frequentist approach to determine the lower bounds on the variance of any unbiased estimator of the quadratic optomechanical coupling. Even without explicitly constructing the estimators, these bounds serve as benchmark values for assessing the performance of those used in experiments. In this context, we analyze the quantum Fisher information (QFI), which is inversely proportional to the lower bound in the Cramér-Rao inequality [12, 22]. We determine the QFI for the optical field state as a function of the unknown strength of the quadratic optomechanical coupling. This is compared with the classical Fisher information (CFI), obtained for balanced-homodyne photodetection measurement scheme. The QFI serves as a measure of the sensitivity of the optical field state to changes in the quadratic optomechanical coupling constant; in other words, it quantifies how much information the quantum state contains about this parameter. However, the CFI assesses the potential to evaluate this sensitivity using a set of classical measurements.
@@ II Model
where $\hat{x}$ is the position operator of the mechanical system, $\hat{p}$ is its momentum operator, $m$ is its mass, and $\Omega$ is the mechanical frequency of motion. The operators $\hat{a},\hat{a}^{\dagger}$ are the photonic ladder operators with the commutation relation $\hat{a},\hat{a}^{\dagger}]=1$, and $\omega(x)$ is the resonant optical frequency of the cavity, which depends on the position $x$ of the mechanical oscillator. If we expand the optical frequency $\omega(x)$ around the equilibrium value $x_{0}$, we obtain
@@
where $\omega^{(n)}(x_{0})$ denotes the $n$th derivative of $\omega(x)$ evaluated at the point $x_{0}$. Here, $g_{1}$ is the linear, and $g_{2}$ is the quadratic optomechanical coupling constant. The parameter $\omega_{c}$ is identified as the frequency of the single mode of the radiation field. It is worth noting that by combining Eqs. (1) and (2), the harmonic potential governing the mechanical system takes on a modified form: $m\Omega^{2}\hat{x}^{2}/2+\hbar g_{2}\hat{a}^{\dagger}\hat{a}\hat{x}^{2}$, where the position is treated as an operator. To preserve the harmonic motion of the mechanical system, we must ensure that for any physically realizable quantum state $\lvert{\Psi}\rangle_{c}$ of the cavity field, the coefficient of $\hat{x}^{2}$ remains positive; that is,
$m\Omega^{2}/2+\hbar g_{2}\,{}_{c}\langle{\Psi}\rvert\hat{a}^{\dagger}\hat{a}\lvert{\Psi}\rangle_{c}>0.$ (3)
This condition not only guarantees harmonic behavior but also serves as a stability criterion for the model, ensuring the system consists of two interacting quantum harmonic oscillators. Under this constraint, the optomechanical Hamiltonian becomes:
@@
It is worth mentioning that the single-mode approach requires the motion of the mechanical oscillator to be adiabatically slow, so the model can ignore the scattering of photons from the mode to other cavity modes [17]. Furthermore, this unitary model assumes that the cavity is lossless and the mechanical oscillator is not subject to any decoherence processes.
@@
where $\hat{I}$ is the identity operator on the Hilbert space of the mechanical oscillator. By setting
$\Omega_{n}=\sqrt{\Omega^{2}+\frac{2\hbar g_{2}n}{m}},$ (7)
which due to (3) is always strictly positive, we can define the ladder mechanical operators $\hat{b}_{n},\hat{b}^{\dagger}_{n}$ as
@@
where
$\omega_{n}=n\omega_{c}+\Omega_{n}/2$ (12)
and $g_{1,n}=ng_{1}\sqrt{\frac{\hbar}{2m\Omega_{n}}}$. By applying the Baker-Campbell-Hausdorff formula, we get [4]
@@
Given an initial state $\hat{\varrho}(0)$, the time evolution of the system is given by the von Neumann equation
$\hat{\varrho}(t)=e^{-i\hat{H}^{\text{quad}}t/\hbar}\hat{\varrho}(0)e^{i\hat{H}^{\text{quad}}t/\hbar}.$ (16)
@@
$\hat{\varrho}(0)=\hat{\rho}(0)\otimes\mathfrak{\hat{m}}(0),$ (17)
where $\hat{\rho}$ and $\mathfrak{\hat{m}}$ are the optical and the mechanical density operator, respectively.
@@
where $d^{2}\alpha=d\mathrm{Re}\{\alpha\}\,d\mathrm{Im}\{\alpha\}$, $\lvert{\alpha}\rangle_{0}$ is a coherent state, and $P(\alpha,\alpha^{*})$ is a Glauber-Sudarshan phase-space distribution [11, 36]. Due to the fact that $\hat{H}^{\text{quad}}$ is block diagonal in the photonic Fock basis, we choose
$\hat{\rho}(0)=\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert$, (19)
@@
$\mathrm{Tr}\{\hat{\rho}(0)\}=1\quad\text{and}\quad\hat{\rho}(0)\geq 0,$ (20)
i.e., $\hat{\rho}(0)$ is a positive semidefinite operator with trace one.
@@
$\hat{\varrho}(t)$ $=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\hat{H}_{n}t/\hbar}\hat{\mathfrak{m}}(0)e^{i\hat{H}_{m}t/\hbar}$ (25)
@@
$\hat{\varrho}(t)$ $=$ $\sum^{\infty}_{n,m=0}a_{n,m}\lvert{n}\rangle_{c}\langle{m}\rvert e^{-i\left[\Phi_{n}(t)-\Phi_{m}(t)\right]}$
@@
The remainder of the calculation will focus solely on the optical density operator, $\hat{\rho}(t)$, at time $t$. This approach is justified by the fact that, in most practical scenarios, the optical field is the only component of the system that is experimentally accessible. The operator $\hat{\rho}(t)$ is obtained by performing a partial trace over the mechanical degrees of freedom, and it can be expressed as
$\hat{\rho}(t)=\mathrm{Tr}_{\text{mech}}\{\hat{\varrho}(t)\}=\sum^{\infty}_{n,m=0}a_{n,m}(t)|n\rangle_{c}\langle m|,$ (31)
@@
The complete expressions for $A$, $b$ and $c$ are provided in Appendix B, all derived from the overlaps of two-phonon coherent states [21], as explicitly given in Eq. (LABEL:eq:A1). As a consistency check, the initial state in Eq. (17) was numerically propagated under $\hat{U}=e^{-i\hat{H}^{\textbf{quad}}t}$(see Eq. (4)) for representative scenarios involving simple parameter values and initial states represented by small size matrices. The mechanical degrees of freedom were then traced out, and agreement with the analytical results in Eq. (36) was observed within numerical accuracy.
The coefficients $a_{n,m}(t)$ of the optical density operator provide a complete description of the interaction between the single mode of the radiation field and the single vibrational mode of the mechanical oscillator, under the assumption that all losses and sources of decoherence are neglected. We now use these results to estimate the strength of the quadratic optomechanical coupling.
@@ III Quantum and Classical Fisher information
$\text{Var}_{g}\geq\frac{1}{N_{\text{meas}}\text{CFI}}\geq\frac{1}{N_{\text{meas}}\text{QFI}},$ (38)
where $g$ is the parameter to be estimated.
@@
In quantum mechanics, the measurement procedure is described by a positive operator-valued measure (POVM), a collection of operators $\{\hat{\Pi}_{i}\}_{i}$, such that any $\hat{\Pi}_{i}\geq 0$ and $\sum_{i}\hat{\Pi}_{i}=\hat{I}$, where $\hat{I}$ is the identity operator on the corresponding Hilbert space. We denote the parameter $g$-dependent optical density operator as $\hat{\rho}_{g}$. The CFI of the quantum state $\hat{\rho}_{g}$ together with the measurements $\hat{\Pi}_{i}$ can be calculated as
$\text{CFI}=\sum_{i}P_{i}(g)\bigg[\frac{\partial\ln P_{i}(g)}{\partial g}\bigg]^{2},$ (39)
where $P_{i}(g)=\mathrm{Tr}\{\hat{\rho}_{g}\hat{\Pi}_{i}\}$ is the probability of getting the outcome $i$ when the true value of the parameter is $g$.
@@
$\text{QFI}=\mathrm{Tr}\{\hat{\rho}_{g}\hat{L}_{g}^{2}\}$ (40)
which is written in terms of the symmetric logarithmic derivative (SLD) $\hat{L}_{g}$ [13], which satisfies the relation
$\frac{\partial\hat{\rho}_{g}}{\partial g}=\frac{1}{2}(\hat{L}_{g}\hat{\rho}_{g}+\hat{\rho}_{g}\hat{L}_{g}).$ (41)
@@
$\hat{L}_{g}=2\int_{0}^{\infty}e^{-\hat{\rho}_{g}x}(\partial_{g}\hat{\rho}_{g})e^{-\hat{\rho}_{g}x}\,dx.$ (42)
We mention here that when the density operator has the property that $\hat{\rho}_{g}^{2}=\hat{\rho}_{g}-h\hat{I}$, with $h\in\mathbb{R}$, which always happens when $\hat{\rho}_{g}$ describes a two-level system, the SLD assumes the simplified form [20]
$\hat{L}_{g}=2\partial_{g}\hat{\rho}_{g}-\frac{1}{2}\partial_{g}\mathcal{P}\hat{\rho}_{g}^{-1},$ (43)
where $\hat{\rho}$ is not a pure state, $\mathcal{P}=\mathrm{Tr}\{\hat{\rho}_{g}^{2}\}$ is the purity of the system, and it is related to $h$ by
$\mathcal{P}=2h+1.$ (44)
This relation can be used to ease the computation of the QFI. In the following we are going to apply Eq. (43) to calculate the QFI for the case when the cavity is populated by a superposition of zero-photon and one-photon states.
In the case of a $2\times 2$ system, we use the balanced homodyne photodetection (BHD) measurements in terms of their projective value measures (PVM). The BHD [39] enables the measurement of the observable
$\hat{X}_{\phi}=\frac{\hat{a}e^{i\phi}+\hat{a}^{\dagger}e^{-i\phi}}{2}$ (45)
with $\phi$ being the phase of the coherent local oscillator. In employing the BHD, we have assumed that the quantum state of the optical field can be accessed without inducing any disturbance.
We rewrite this to account for the two-level system scenario that we are considering by substituting the operators $\hat{a},\hat{a}^{\dagger}$ with the $2\times 2$ counterparts $\hat{\sigma}^{-},\hat{\sigma}^{+}$, respectively. The corresponding PVM $\{\hat{\Pi}_{X_{\phi}}\}$ is given by the projectors on the two eigenstates of $\hat{X}_{\phi}$,
$\hat{\Pi}_{X_{\phi}}(k)=|e_{k}\rangle\langle e_{k}|,\quad k=1,2.$ (46)
Figure 1: Wigner phase-space distribution of the optical field initially prepared in an coherent state, with $\alpha=10$ and system parameters defined by Eq. 47. In the figures, $g_{1}$ takes the values from 0, 11 and $10$ left to right, while $g_{2}$ top to bottom. The plots are snapshots of the distribution taken at time equals $T_{\omega}$, defined in Eq. (49). With the border outlined in red, the top left figure is both a snapshot of the system at $T_{\omega}$ for $g_{1}=g_{2}=0$ but also the initial state for all the other figures at $t=0$.